show the equation used for the solution.
The length of a rectangle is 2 in. more than twice its width. If the perimeter of the rectangle is 34 in., find the dimensions of the rectangle.

Answer by scott8148(6628)

L=2W+2 ... P=2L+2W ... 34=2(2W+2)+2W ... 34=4W+4+2W ... 30=6W ... 5=W ... L=12